# Collaborative motion planning for multi-manipulator systems through Reinforcement Learning and Dynamic Movement Primitives

### v1

@@ I INTRODUCTION
Figure 1: Combining RL with enhanced DMP, the proposed method generates online trajectories in real-time to enhance collaboration.
@@ III-B2 Collaborative Execution
$d=\begin{cases}||\bm{x}-\bm{x}^{\prime}||_{2},&\text{if }||\bm{x}-\bm{x}^{\prime}||_{2}\leq r_{s}\\ 0,&\text{otherwise}\end{cases}$ (22)
$r_{s}$ is a user defined parameter computed based on the dimensions of the end-effector to prevent collision similar to an inflation radius. Since the relaxation of the execution is imparted using the parameter $\alpha_{s}$ the exponential stability of the trajectory is remains intact.
@@ IV EXPERIMENTS AND VALIDATION
To validate the effectiveness of the proposed method we designed multiple experiments. We use two UR5e robotic manipulators with Robotiq 2f-85 for end-effector on each arm. We refer to them as arm-1 and arm-2 for the remainder of the section. The experimental setup is based in the PyBullet simulation environment. We study the ability of the proposed method to i) devise online collision-free trajectories, ii) prevent collision amongst the arms and iii) complete distinct tasks collaboratively. For the purpose of simplicity we assume complete observability, i.e. the position and the velocity of each element in the scene is known and can be measured in real time.
@@
Figure 3: Setup in PyBullet environment for crossing arms. Top row shows the trajectory for the case with the proposed ONColDMP, whereas bottom row shows traditional DMP without collaborative term.
Figure 4: Comparison of the DMP trajectories for end-effector with the improvised ONCol DMP (top row) v/s without collaborative phase control term (bottom row).
@@ IV-A Cross-Trajectory Collaborative Task
We study the improvement of the trajectory execution as proposed in Sec.III-B2. To better highlight the strength of the method we purposely devise trajectories with starting and ending goal positions which require the two arms to cross each other. As shown in Fig. 4, introducing the collaborative reduces the execution speed of the second arm while the first arm continues. This leads to a much smaller deviation in the trajectory as compared to a case where the other arm is only considered as an obstacle. For the purpose of this experiment $\hat{\alpha}=25/3$ and $r_{s}=0.25$. All the other parameters were the same for the two trajectories. The maximum deviation for arm-1 and arm-2 were 0.01m for the first arm and 0.13m for the second arm respectively in case of ONCol-DMP. Without any change in the phase control variable, the maximum deviation of the first arm was measured at 0.12m and for the second arm it was measured at 0.14m respectively. Since, only the second robot was equipped with the enhanced $\alpha_{s}$, the reduced deviation for arm-1 coincides with the hypothesis.
@@
Figure 6: PyBullet setup showing collaborative stacking of blocks with two arms.
@@ IV-B Long Sequence Collaborative Tasks
Lastly, the proposed method was used to complete three long sequence tasks, namely block stacking, water transfer and cleaning a table [^3] For videos and more details of different tasks please visit the project website. The task constraints are provided as a sequence of critical-configurations as discussed in Sec. III-A. The proposed method lead to successful collision free completion of task for all three cases. Figure 5 shows the trajectory for the block stacking task involving two UR5e arms. Since the stack has to be made at a single place, when arm-1 places a block, it can act as an obstacle for arm-2. However, as evident from the plots, the motion along the y-axis, starts slowing down for arm-2 after the 1.6s mark, whereas the arm-1 continues the motion at the desired speed. Once arm-1 places the block, it moves out of the way and arm-2 can continue placing block. The delay causes the arm-2 DMP to be executed for longer duration.
@@ V CONCLUSIONS AND FUTURE WORK
Through various experiments we identified that the proposed method can: i) generate trajectories for multiple arms given their respective task constraints, ii) avoid obstacles dynamically without parameter re-tuning, and iii) accomplish collaborative tasks by heuristic collaborative phase control. The proposed technique can be deployed to generate new trajectories online and in real time. Users can enhance the ability of the robots to generate contextual trajectories by either adding more demonstrations or modifying the existing ones. For example, in a robotic assembly task the demonstrated skills can specifically focus on screwing/unscrewing, bolting, stacking etc., giving the users a more flexible setup. The proposed method works as an off-the-shelf plug and play tool since it only requires training a single RL agent and then deploying to multiple arms without re-training facilitating collaboration. The independence of the trained RL agent also allows user to impart distinct behaviors/roles to different arms. Although the proposed method was validated using two manipulators, it can also be extended to a larger number of agents.

### v2

@@ III-B2 Collaborative Execution
$d=\begin{cases}||\bm{x}-\bm{x}^{\prime}||_{2},&\text{if }||\bm{x}-\bm{x}^{\prime}||_{2}\leq\epsilon_{s}\\ \infty,&\text{otherwise}\end{cases}$ (22)
$\epsilon_{s}$ is a user defined parameter computed based on the dimensions of the end-effector to prevent collision similar to an inflation radius. Since the relaxation of the execution is imparted using the parameter $\alpha_{s}$ the exponential stability of the trajectory is remains intact.
@@ IV EXPERIMENTS AND VALIDATION
To validate the effectiveness of the proposed method we designed multiple experiments in simulation and on hardware. For the simulation purpose we use two UR5e robotic manipulators with Robotiq 2f-85 for end-effector on each arm. We refer to them as arm-1 and arm-2 for the remainder of the section. The simulation experimental setup is based in the PyBullet simulation environment. For the hardware setup we use a UR5e (6 DoF) manipulator with OnRobot RG2 gripper and a Kinova Gen-3 (7 DoF) manipulator with Robotiq 2f-85 gripper as shown in Fig. 2.
We study the ability of the proposed method to i) devise online collision-free trajectories, ii) prevent collision amongst the arms and iii) complete distinct tasks collaboratively. For the purpose of simplicity we assume complete observability, i.e. the position and the velocity of each element in the scene is known and can be measured in real time.
@@
Figure 2: Hardware Implementation of ONCol DMP on a UR5e and Kinova Gen-3 robotic arm.
Figure 3: Setup in PyBullet environment for crossing arms. Top row shows the trajectory for the case with the proposed ONColDMP, whereas bottom row shows traditional DMP without collaborative phase control.
Figure 4: Comparison of the DMP trajectories in cartesian space for end-effector with the improvised ONCol DMP (top row) v/s without collaborative phase control term (bottom row).
Figure 5: PyBullet setup showing collaborative with two arms.(a-c) Block stacking, (d-f) Table cleaning, and (g-i) water transfer.
@@ IV-A Cross-Trajectory Collaborative Task
We study the improvement of the trajectory execution as proposed in Sec.III-B2. To better highlight the strength of the method we purposely devise trajectories with starting and ending goal positions which require the two arms to cross each other. We simulate the experiment in the PyBullet environment as shown in Fig. 3. As indicated in Fig. 4, introducing the collaborative behavior reduces the execution speed of arm-2 while the arm-1 continues to execute its trajectory. This leads to a much smaller deviation in the trajectory as compared to a case where the other arm is only considered as an obstacle. For the purpose of this experiment $\hat{\alpha}=25/3$ and $r_{s}=0.25$. All the other parameters were the same for the two trajectories. Table I shows the comparison of performance for the two cases. We observe the maximum deviation a parameter of performance. It is evident that utilizing the collaborative term reduces the deviation in the motion by introducing a sequential-like approach. Since, only the second robot was equipped with the enhanced $\alpha_{s}$, the reduced deviation for arm-1 coincides with the hypothesis.
TABLE I: Max. Deviation of Arms in Crossing Task
Arm-1 Arm-2
DMP 0.12 m 0.14 m
ONCol DMP 0.01 m 0.13 m
@@ IV-B Long Sequence Collaborative Tasks
We further extend our method to demonstrate the ability to handle distinct long sequence tasks as shown in Fig. 5, namely block stacking Fig. 5 (a-c), table cleaning Fig. 5 (d-f), and water pouring Fig. 5 (g-i). The task constraints are provided as a sequence of critical-configurations as discussed in Sec. III-A. The proposed method lead to successful collision free completion of task for all three cases. Due to reason of space constraints we showcase the behavior of the block stacking task  [^3] For videos and more details of different tasks please visit the project website: https://sites.google.com/virginia.edu/oncoldmp/home. Figure 6 shows the trajectory for the block stacking task involving two UR5e arms. Since the stack has to be made at a single place, when arm-1 places a block, it can act as an obstacle for arm-2. However, as evident from the plots, the motion along the y-axis, starts slowing down for arm-2 after the 1.6 s mark, whereas the arm-1 continues the motion at the desired speed. Once arm-1 places the block, it moves out of the way and arm-2 can continue placing block. The delay causes the arm-2 DMP to be executed for longer duration.
We further validated our technique on hardware setup as shown in Fig. 2 using Kinova 7-DoF arm and a UR5e 6-DoF arm further highlighting the adaptability of the proposed method.
IV-C Multi-Arm Scenario
Since the proposed method is not limited by the number of agents and the training of LfD agents is independent, this allows us to extend the method without any changes to the larger number of arms. We validate this by extending our method to a multi-arm setup as shown in Fig. 7 for a block stacking task.
Figure 7: PyBullet setup showing collaborative task of stacking with three UR5e arms.
@@ V CONCLUSIONS AND FUTURE WORK
Through various experiments we identified that the proposed method can: i) generate trajectories for multiple arms given their respective task constraints, ii) avoid obstacles dynamically without parameter re-tuning, and iii) accomplish collaborative tasks by heuristic collaborative phase control. The proposed technique can be deployed to generate new trajectories online and in real time. Users can enhance the ability of the robots to generate contextual trajectories by either adding more demonstrations or modifying the existing ones. For example, in a robotic assembly task the demonstrated skills can specifically focus on screwing/unscrewing, bolting, stacking etc., giving the users a more flexible setup. The proposed method works as an off-the-shelf plug and play tool since it only requires training a single RL agent and then deploying to multiple arms without re-training facilitating collaboration. The independence of the trained RL agent also allows user to impart distinct behaviors/roles to different arms.